if the length of a rectangle is twice the width and the permiter is 120 centimeters what are the lengths and widths?
Answer by emargo19(101)

Let W=x
let L=2x
2(l+2x)=120
2x+4x=120
6x=120
x=20
w=20
l=(20)(2)=40
Length=40cm
Width=20cm